The increase in the volume of 10cm cube of mercury when the temperature rise by 100 degree census 150.82cm what is the cubic expansivity of mercury

1
Expert's answer
2021-04-14T20:59:58-0400

By definition, the cubic expansivity is:


"\alpha_V = \dfrac{1}{V}\dfrac{\Delta V}{\Delta T}"

where "V = 10cm^3" is the initial volume, "\Delta V = 150.82cm^3" is the increase in the volume, "\Delta T = 100\degree C = 100K" is the increase in temperature. Thus, obtain:


"\alpha_V = \dfrac{1}{10cm^3}\dfrac{150.82 cm^3}{100K} \approx 0.15K^{-1}"

Answer. "0.15\space K^{-1}".